# Learning-Based Two-Way Communications: Algorithmic Framework and Comparative Analysis

### v1

@@ Abstract
Machine learning (ML)-based feedback channel coding has garnered significant research interest in the past few years. However, there has been limited research exploring ML approaches in the so-called “two-way” setting where two users jointly encode messages and feedback for each other over a shared channel. In this work, we present a general architecture for ML-based two-way feedback coding, and show how several popular one-way schemes can be converted to the two-way setting through our framework. We compare such schemes against their one-way counterparts, revealing error-rate benefits of ML-based two-way coding in certain signal-to-noise ratio (SNR) regimes. We then analyze the tradeoffs between error performance and computational overhead for three state-of-the-art neural network coding models instantiated in the two-way paradigm.
@@
Channel coding, two-way communications, AI/ML-based communications, feedback coding
@@ I Introduction
In modern wireless communications systems, bidirectional exchange of information between users can be critical, as in, for example, communications between a base station and a satellite. Such a system is referred to as a “two-way channel” [1]. In particular, focus has been dedicated to systems in which additive white Gaussian noise (AWGN) is added independently in both directions, called Gaussian two-way channels (GTWCs). A key early result on two-way coding demonstrated that when noise is independently introduced in both directions, the optimal channel capacity is attained by treating each direction as a separate one-way channel [2]. Additionally, [3] established potential benefit for error exponents in cooperative GTWCs over non-cooperative cases.
Two-way feedback coding is itself an extension of the Gaussian one-way channel model (GOWC), which began with the pioneering work by Schalkwijk and Kailath in [4]. While this one-way scheme was effective in cases of no feedback noise, it would be several decades before the development of GOWC schemes which were effective with noisy feedback [5], with more recent advances in classical (linear) methods coming with DeepPolar in [6]. More recently, works on GOWCs began employing machine learning (ML) to perform the encoding and decoding tasks, which, as discussed below, has led to orders of magnitude improvement in performance metrics in certain signal-to-noise ratio (SNR) regimes.
@@ I-A Non-Linear One-way Coding
Another major advance came from LightCode (LC) [12], which found that in the right use-cases, the attention mechanisms in GBAF provided at best marginal benefit in system performance. Building upon GBAF’s notion of a feature extractor, [12] formulated a system with only a small feed-forward network with a skip connection. LightCode’s encoder/decoder architecture achieved comparable performance to GBAF with only 10% of the trainable weights. In reducing the number of weights, LC is more time- and space-efficient than its transformer- and RNN-based counterparts.
Lastly, the authors of [13] developed an RNN-based model, RobustCode, featuring an attention module at its decoder’s output. RobustCode demonstrated SOTA performance and is the basis for the two-way RNN-based scheme discussed below.
@@ I-B Two-way Coding Schemes
Despite there being considerable progress in the development of one-way methodologies, as well as some recent explorations of broadcast channels [14, 15], there exists a relative dearth of work examining feedback coding in two-way systems. Recently, [16] was the first to attempt to explore finite block-length two-way coding, presenting both a linear system developed akin to [5] and an extension of RobustCode to the two-way case (TWRNN) which is trained to minimize the sum-BLER of the users. The work tackles what it identifies as the three primary difficulties of GTWCs: (1) the coupling of the two users’ encoding processes, (2) the need for decoders capable of parsing through this coupling, and (3) the need for effective power allocation within a given power constraint. TWRNN displays robustness by performing well in low SNR regimes, while the linear scheme excels at high SNR.
@@ I-C Summary of Contributions
Our contributions are summarized as follows:
  1. 1.
    We present a general framework for optimizing sum block error rate in non-linear encoder/decoder systems. With this model, we develop new two-way architectures, Two-way Block Attention Feedback (TWBAF) and Two-way LightCode (TWLC), based on the one-way models found in [9] and [12], respectively.
  2. 2.
    We assess the tradeoff between one-way and two-way coding by comparing TWLC against a new active feedback-enabled variant which we develop, Active Feedback LightCode (ALC). Therein, we find distinct regions over which one- and two-way systems demonstrate relatively superior performance.
  3. 3.
    We perform simulations across different SNR regimes to investigate the performance tradeoffs between the models in terms of BLER and computational overhead using floating point operations (FLOPS).
II System Model
We consider a system with two users, $i=1,2$, each aiming to transmit, respectively, bitstreams (messages) $\bm{b}_{i}\in\{0,1\}^{K_{i}}$ at rate $R$ under the two-way feedback coding paradigm. Let $t=1,2,\dots,T$ be the channel use indices over which the bitstreams are sent, where the symbol $c_{i}^{t}\in\mathbb{R}$ is the symbol sent by user $i$ at time $t$, as generated by a transmit-side neural network-based encoder; each user also possesses a receive-side ML-enabled decoder for processing received symbols. The received symbol at user $i$ is then
@@ II System Model
where $n_{j}^{t}\sim\mathcal{N}(0,\sigma_{j}^{2})$ is independent and identically distributed AWGN $\forall t$. We collect these as $\bm{n}_{i}=[n_{i}^{1},n_{i}^{2},\dots,n_{i}^{T}]^{\intercal}$ and $\bm{y}_{i}=[y_{i}^{1},y_{i}^{2},\dots,y_{i}^{T}]^{\intercal}$ and subject each user’s transmissions to a power constraint given by
$\mathbb{E}_{\bm{b}_{i},\bm{n}_{1},\bm{n}_{2}}\left[\sum_{t=1}^{T}c_{i}^{t}\right]\leq P_{i}T,$ (2)
@@
III Non-Linear Optimization Scheme and Encoder/Decoder Architectures
We begin by presenting the methods by which BLER is jointly minimized across the users. We then introduce several new specific architectures for use in subsequent simulations.
III-A Model Inputs and Outputs
Similar to [9], we define transmit-side knowledge vectors $\bm{q}_{i}^{t}\in\mathbb{R}^{K_{i}+2(T-1)}$ as the inputs to the encoder. Define the overall functional representation of generating transmit symbols at the encoder at time $t$ as $c_{i}^{t}=E_{i}(\bm{q}_{i}^{t})$, which is power-constrained under the trainable power reallocation scheme presented in [16]. At the beginning of each transmission block, the knowledge vectors are instantiated for channel use 1 as follows:
@@ III-A Model Inputs and Outputs
$\bm{0}_{2(T-1)}$ denotes a length-${2(T-1)}$ padding vector which will be replaced by previously sent symbols, $c_{i}^{t^{\prime}<t}$, and by previously received information $\tilde{y}_{i}^{t^{\prime}<t}$, which equals either $y_{i}^{t^{\prime}<t}$ or $y_{i}^{t^{\prime}<t}-c_{i}^{t^{\prime}<t}$, over subsequent channel uses [^1] For simplicity of notation, we henceforth suppress the tilde over $y$.. At $t=2$, the knowledge vectors are updated as
@@
with the process continuing iteratively until the final channel use at time $T$:
@@
After $T$ channel uses, the decoder at user $i$ takes receive-side knowledge vector, $\bm{r}_{i}\in\mathbb{R}^{K_{i}+T}$, as input, defined as $\bm{r}_{i}=[\bm{b}_{i},\bm{y}_{i}]^{\intercal}$. Let $D_{i}=\mathcal{O}_{i}\circ\widehat{D}_{i}$ denote the functional representation of user $i$’s decoder. [^2] Decoder $D_{i}$ is housed at user $j\neq i$. $\widehat{D}_{i}$ is the output of the receive-side neural network, and $\mathcal{O}_{i}$ is given shortly. Let $\widehat{\bm{d}}_{i}=\widehat{D}_{i}(\bm{r}_{j})\in\Delta^{2^{K_{i}}},i\neq j$, where $\Delta^{2^{K_{i}}}$ is the standard $2^{K_{i}}$-simplex. Let $\mathcal{O}_{i}:\bm{k}\rightarrow\{0,1\}^{K_{i}}$ be a bijection mapping $\bm{k}$, a one-hot vector whose non-zero entry is at index $k$, to a unique length-$K_{i}$ binary bitstream. Thus, the estimated bitstream for user $i$ can be obtained as
$\widehat{\bm{b}}_{i}=\mathcal{O}_{i}(\operatorname*{argmax}{\widehat{\bm{d}}_{i}}).$ (6)
@@
III-B Optimization of Encoders and Decoders
To optimize the transmission and decoding of the bitstreams, we consider the natural metric of minimizing error probability. The principle metric we employ is that of block error rate (BLER), which is the ratio of the number of incorrect estimated bitstreams to the total number of bitstreams sent:
@@ III-B Optimization of Encoders and Decoders
The goal of the system is therefore to minimize the sum BLER of the users by jointly optimizing over their encoders and decoders, subject to (2). Writing the BLERs as functions of the relevant encoder and decoder, we are left with the following optimization problem
@@
Since $\bm{d}_{i}$ is mapped to a unique sequence of ones and zeros via $\mathcal{O}_{i}$, we have that
$\mathsf{Pr}\left[\{\bm{b}_{i}\neq\widehat{\bm{b}}_{i}\}\right]=\mathsf{Pr}\left[\{\mathcal{O}_{i}^{-1}\left(\bm{b}_{i}\right)\neq\widehat{\bm{d}}_{i}\}\right],~{}i=1,2.$ (9)
The optimization problem presented in (8) can therefore be considered as a multi-class classification problem over $2^{K_{i}}$ classes at the decoders’ outputs, for which we adopt the standard metric of cross-entropy loss. Let $\bm{d}_{i}=\mathcal{O}_{i}^{-1}\left(\bm{b}_{i}\right),~{}i=1,2$, and let the models for each user be parameterized by $\theta_{i}$, respectively. The per-user and overall loss are then
$\mathcal{L}_{\theta_{i}}(\{\bm{d}_{i},\widehat{\bm{d}}_{i}\})=-\sum_{\ell=1}^{2^{K}_{i}}d_{i}[\ell]\log\left(\widehat{d}_{i}[\ell]\right),$ (10)
$\textrm{and}~{}~{}~{}\mathcal{L}=\mathcal{L}_{\theta_{1}}+\mathcal{L}_{\theta_{2}}.~{}~{}~{}~{}~{}~{}~{}~{}$ (11)
@@
denote the function which updates the transmit-side knowledge vectors using (3)-(5). The overall GTWC training methodology outlined above is summarized in Alg. 1. While we demonstrate the training cycle for a single bitstream, we note here that a full model should be trained for a requisite number of epochs using batched inputs of knowledge vectors.
III-C Channel Coding Architectures
Using the general framework developed above, we present three new channel coding architectures which will function as the transmit- and receive-side neural networks: a one-way variant of LC leveraging active feedback and two two-way systems based on LC and GBAF. Block diagrams for LC and GBAF encoders/decoders can be found in Fig. 2.
Figure 2: Block diagrams for GBAF and LC encoder/decoder neural networks. The feature extractors (FE) and multi-layer perceptrons (MLPs) are fully-connected feed-forward networks; LC contains a skip-connection which connects its FE’s first layer to the FE’s last last. The positional encoding and transformer (Xformer) encoder block follow standard implementations as discussed in [8]. Model layer specifications for TWLC and TWBAF follow roughly those outlined in [12, 9], with modifications as indicated in Sec. III-C.
Two-Way LightCode-Based System: We begin with the overall model shape described in [12] (Figs. 12 & 13 therein). In adapting it to a two-way system, we double the number of neurons per hidden layer with the goal of making the model more competitive in the BLER metric. Additionally, layer normalization is added after the last layer of the feature extractor. We call this system Two-Way LightCode (TWLC).
One-Way Active Feedback LightCode System: To facilitate a fair comparison between one-way and two-way models under the assumption that both transmission directions can utilize an ML-based encoder, we develop Active Feedback LightCode (ALC) [^3] We constrain the transmit power of the feedback encoder so that it matches the power constraint for the encoder transmitting the message, i.e., if $P=1$ and $T=9$, the RHS of (2) is $9$ for both the forward and feedback encoders.. To implement ALC, we use the TWLC encoder, with User 2 always fed $\bm{b}_{2}=\bm{0}$ as a dummy initialization; the decoder is only fed $\bm{y}_{2}$ since $\bm{b}_{2}$ provides no pertinent information. The overall loss is given by User 1’s CE loss. The layer dimensions of ALC in Tab. II are matched exactly to those in Figs. 12 & 13 in [12] for consistent comparison; elsewhere, its dimensions are matched to those of TWLC.
GBAF-Based System: Beginning with the encoder/decoder modules presented in [9], we insert, with some abuse of notation, the function $\tanh{(\bm{X}-\mathbb{E}[\bm{X}])}$, where $\bm{X}\in\mathbb{R}^{B\cdot\ell}$, $B$ is the batch size, $\ell$ is the number of rows in the input vector, and $\mathbb{E}[\bm{X}]\in\mathbb{R}$ is over all the elements in the batch. Additionally, bits are not BPSK modulated at the model’s input. Without these updates, the two-way model experienced training instability, failing to converge in all cases of $SNR_{1}=-1$ dB. We employ feature extractor C with ReLU activations without any modification, while also matching the original specifications of the dimensions of the encoder/decoder attention modules. We refer to this specific implementation as Two-way Block Attention Feedback Coding (TWBAF).
1 Input: Bitstream length $K_{i}$, channel uses $T$, noise power $\sigma_{i}^{2}$, power constraint $P_{i}$, model parameters $\theta_{i}$ (for $i=1,2$).
2 Generate bitstream: $\bm{b}_{i}\in\{0,1\}^{K_{i}}$.
@@ 3 Make knowledge vectors: $\bm{q}_{i}^{1}\leftarrow[\bm{b}_{i},\bm{0}^{2(T-1)}]$
5       Generate transmit symbols, ${c}_{i}^{t}\leftarrow E_{i}(\bm{q}_{i}^{t})$, then transmit to receiver.
@@
9Make received knowledge vector: $\bm{r}_{i}\leftarrow[\bm{b}_{i},\bm{y}_{i}]$
10 Decode symbols: $\widehat{\bm{d}}_{j}\leftarrow\widehat{D}_{i}(\bm{r}_{i})$.
@@
12Back-propagate. Update gradient scheduler.

### v2

@@ Abstract
Machine learning (ML)-based feedback channel coding has garnered significant research interest in the past few years. However, there has been limited research exploring ML approaches in the so-called “two-way” setting where two users jointly encode messages and feedback over a shared channel. In this work, we present a general architecture for ML-based two-way feedback coding, and show how several popular one-way schemes can be converted to the two-way setting through our algorithmic framework. We compare such schemes against one-way counterparts, revealing error-rate benefits of ML-based two-way coding in certain signal-to-noise ratio (SNR) regimes. We then analyze the tradeoffs between error performance and computational overhead for three state-of-the-art neural network coding models instantiated in the two-way paradigm.
@@
Channel coding, AI/ML for communications
@@ I Introduction
In modern communication systems, bidirectional exchange of information can be critical, as in, for example, communications between a base station and a satellite. Such a system is referred to as a “two-way channel” [1]. Focus has been dedicated to settings in which additive white Gaussian noise (AWGN) is added independently in both directions, called Gaussian two-way channels (GTWCs). A key early result on two-way coding demonstrated that when noise is independently introduced in both directions, the optimal channel capacity is attained by treating each direction as a separate one-way channel [2]. More recently, [3] established potential benefit for error exponents in cooperative GTWCs over non-cooperative cases.
Two-way (TW) feedback coding is itself an extension of the Gaussian one-way (OW) feedback channel model (GOWC), pioneered in [4]. While this initial OW scheme was effective in cases of no feedback noise, it would be several decades before the development of GOWC schemes which were effective with noisy feedback [5, 6]. More recently, works on GOWCs began employing machine learning (ML) to perform the encoding and decoding tasks. As discussed below, these ML techniques have led to significant improvement in performance metrics in certain signal-to-noise ratio regimes.
@@ I-A Non-Linear One-way Coding
Another major advance came from LightCode (LC) [12], which found that in certain use-cases, the attention mechanisms in GBAF provide marginal benefit in system performance. Building upon GBAF’s notion of a feature extractor, [12] formulated a system with only a small feed-forward network with a skip connection. LightCode’s encoder/decoder architecture achieved comparable performance to GBAF with only 10% of the trainable weights. In reducing the number of weights, LC is more time- and space-efficient than its transformer- and RNN-based counterparts. Lastly, the authors of [13] developed an RNN-based model, RobustCode, featuring an attention module at its decoder’s output. RobustCode is the basis for the two-way RNN-based scheme discussed below.
@@ I-B Two-way Coding Schemes
Despite considerable progress in the development of one-way codes, as well as some recent explorations of broadcast channels [14, 15], there exists relatively little work examining feedback coding in two-way systems. Recently, [16] was the first to explore finite block-length two-way coding, presenting both a linear system developed akin to [5] and an extension of RobustCode to the two-way case (TWRNN) which is trained to minimize the sum-BLER of the users. The work tackles three difficulties of GTWCs: (1) the coupling of the two users’ encoding processes, (2) the need for decoders capable of parsing through this coupling, and (3) the need for effective power allocation within a given power constraint. TWRNN displays robustness by performing well in low SNR regimes, while the linear scheme excels at high SNR.
@@ I-C Summary of Contributions
  • We present an algorithmic framework for optimizing sum block error rate in non-linear encoder/decoder systems with a generalized input space which allows the use of more of the available past information. With this architecture, we develop new ML-based two-way coding schemes, Two-way Block Attention Feedback (TWBAF) and Two-way LightCode (TWLC), based on the one-way models found in [9, 12].
  • We assess the tradeoffs between OW and TW coding schemes by comparing (i) TWLC against a novel active feedback-enabled OW variant which we develop, Active LightCode, and (ii) the linear two-way scheme of [16] against the inner code of [5]. Results show distinct improvements can be wrought from the TW paradigm, though this is both SNR-regime- and model-dependent.
  • We investigate the performance-complexity tradeoffs between different ML-based two-way codes through numerical experiments comparing BLER and analysis of floating point operations (FLOPS) executed by each. In doing so, we provide insights into the efficacy of different model types for future works in two-way system design.
II System Model and Optimization
We consider two users, $i=1,2$, each aiming to transmit bitstreams $\bm{b}_{i}\in\{0,1\}^{K_{i}}$ at rate $R_{i}=K_{i}/T$. Let $t=1,\dots,T$ be the channel use indices, and let $c_{i}^{t}\in\mathbb{R}$ be the symbol sent by user $i$ at time $t$, as generated by a transmit-side neural network-based encoder; each user also possesses a receive-side ML-enabled decoder. The received symbol at user $i$ is then
@@ II System Model and Optimization
where $n_{j}^{t}\sim\mathcal{N}(0,\sigma_{j}^{2})$ is i.i.d. $\forall t$. We collect these as $\bm{n}_{i}=[n_{i}^{1},\dots,n_{i}^{T}]$, $\bm{c}_{i}=[c_{i}^{1},\dots,c_{i}^{T}]$, and $\bm{y}_{i}=[y_{i}^{1},\dots,y_{i}^{T}]$ and subject each user to a power constraint given by
$\mathbb{E}_{\bm{b}_{i},\bm{n}_{1},\bm{n}_{2}}\left[\sum\nolimits_{t=1}^{T}c_{i}^{t}\right]\leq P_{i}T,$ (2)
@@
II-A ML Model Inputs and Outputs
Similar to [9], we define transmit-side knowledge vectors $\bm{q}_{i}^{t}\in\mathbb{R}^{K_{i}+2(T-1)}$ as the encoders’ inputs. Define the overall functional representation of generating transmit symbols at the encoder at time $t$ as $c_{i}^{t}=E_{i}(\bm{q}_{i}^{t})$, [^1] In [1], Shannon considers a sequence of time-indexed encoding functions. It is more convenient for us to consider a single encoding function which is implicitly time-indexed by the knowledge vectors to simplify the presentation of the ML-based optimization scheme. which is power-constrained under the trainable power reallocation scheme from [16]. The knowledge vectors are instantiated at $t=1$ as:
@@ II-A ML Model Inputs and Outputs
$\bm{0}_{2(T-1)}$ denotes a length-${2(T-1)}$ padding vector that will be replaced by previously sent symbols, $c_{i}^{t^{\prime}<t}$, and by previously received information $\tilde{y}_{i}^{t^{\prime}<t}$, which equals either $y_{i}^{t^{\prime}<t}$ or $y_{i}^{t^{\prime}<t}-c_{i}^{t^{\prime}<t}$, over subsequent channel uses. [^2] For simplicity of notation, we henceforth suppress the tilde over $y$. At $t=2$, the knowledge vectors are updated as
@@
continuing iteratively until the final channel use at time $T$:
@@
After $T$ channel uses, we form receive-side knowledge vectors as $\bm{r}_{i}=[\bm{b}_{i},\bm{y}_{i},{\color[rgb]{0,0,0}\bm{c}_{i}}]^{\intercal}$. Let $D_{i}=\mathbf{O}_{i}\circ\operatorname*{argmax}(\widehat{D}_{i})$ denote the functional representation of user $i$’s decoder. [^3] Decoder $D_{i}$ is housed at user $j\neq i$. $\widehat{D}_{i}$ is the output of the receive-side neural network, and $\mathbf{O}_{i}$ is given shortly. Let $\widehat{\bm{d}}_{i}=\widehat{D}_{i}(\bm{r}_{j})\in\Delta^{2^{K_{i}}},i\neq j$, where $\Delta$ is the unit-simplex. Let $\mathbf{O}_{i}:\bm{k}\rightarrow\{0,1\}^{K_{i}}$ be a bijection mapping $\bm{k}$, a one-hot vector with non-zero entry at index $k$, to a length-$K_{i}$ binary bitstream. The estimated bitstream for user $i$ is then
$\widehat{\bm{b}}_{i}=\mathbf{O}_{i}(\operatorname*{argmax}{\widehat{\bm{d}}_{i}}).$ (6)
@@
II-B Optimization of Encoders and Decoders
We consider the metric of minimizing block error rate (BLER), which is the ratio of the number of incorrect estimated bitstreams to the total number of bitstreams sent:
@@ II-B Optimization of Encoders and Decoders
The goal of the system is therefore to minimize the sum BLER of the users by jointly optimizing over their encoders and decoders, subject to (2), i.e.,
@@
Since $\bm{d}_{i}$ is mapped to a unique sequence of ones and zeros via $\mathbf{O}_{i}$, we have that
$\mathsf{Pr}\left[\{\bm{b}_{i}\neq\widehat{\bm{b}}_{i}\}\right]=\mathsf{Pr}\left[\{\mathbf{O}_{i}^{-1}\left(\bm{b}_{i}\right)\neq\widehat{\bm{d}}_{i}\}\right],~{}i=1,2.$ (9)
The optimization problem in (8) can therefore be considered a classification problem over $2^{K_{i}}$ classes, for which we adopt cross-entropy (CE) loss. Let $\bm{d}_{i}=\mathbf{O}_{i}^{-1}\left(\bm{b}_{i}\right)$, and parameterize user $i$’s models by $\theta_{i}$. The per-user and overall loss are then
$\mathcal{L}_{\theta_{i}}(\{\bm{d}_{i},\widehat{\bm{d}}_{i}\})=-\sum\nolimits_{\ell=1}^{2^{K_{i}}}d_{i}[\ell]\log\left(\widehat{d}_{i}[\ell]\right)$ (10)
$\text{and}~{}\mathcal{L}=\mathcal{L}_{\theta_{1}}+\mathcal{L}_{\theta_{2}}.$ (11)
@@
denote the function which updates the transmit-side knowledge vectors using (3)-(5). The overall GTWC optimization scheme is summarized for a single bitstream in Alg. 1.
III Channel Coding Architectures
Using the framework developed above, we present three new ML-based channel coding architectures: a OW variant of LC leveraging active feedback and two TW systems based on LC and GBAF. Block diagrams for LC and GBAF encoders/decoders can be found in Fig. 2.
Figure 2: Block diagrams for GBAF and LC. The feature extractors (FE) and multi-layer perceptrons (MLPs) are fully-connected feed-forward networks; LC contains a skip-connection from its FE’s first layer to its last. The positional encoding and transformer (Xformer) encoder block follow the standard implementations in [8]. Model layer specifications for TWLC and TWBAF follow roughly those outlined in [12, 9], with modifications as indicated in Sec. III.
TW LightCode-Based System (TWLC): We begin with the overall model shape described in [12] (Figs. 12 & 13 therein). In adapting it to a two-way system, we double the number of neurons per hidden layer with the goal of making the model more competitive in the BLER metric while keeping the model relatively small as intended. Additionally, layer normalization is added after the last layer of the feature extractor for stability. While TWLC was bested in BLER by some other models (Fig. 3), its training was by far the most robust, experiencing neither power-allocation issues nor training instability seen by TWRNN and TWBAF (below). Its smaller size also makes it more amenable to operation on smaller devices.
1 Input: Bitstream length $K_{i}$, block-length $T$, noise power $\sigma_{i}^{2}$, power constraint $P_{i}$, model parameters $\theta_{i}$ (for $i=1,2$).
2 Generate bitstreams: $\bm{b}_{i}\in\{0,1\}^{K_{i}}$.
@@ 3 Make knowledge vectors: $\bm{q}_{i}^{1}\leftarrow[\bm{b}_{i},\bm{0}^{2(T-1)}]$
5       Make transmit symbols, ${c}_{i}^{t}\leftarrow E_{i}(\bm{q}_{i}^{t})$, then transmit.
@@
9Make received knowledge vector: $\bm{r}_{i}\leftarrow[\bm{b}_{i},\bm{y}_{i},\bm{c}_{i}]$
10 Decode symbols: $\widehat{\bm{d}}_{j}\leftarrow\widehat{D}_{j}(\bm{r}_{i})$.
@@
12Back-propagate. Update gradient scheduler (if needed).
@@
OW Active Feedback LightCode System (ALC): To facilitate a fair comparison between OW and TW models, assuming that both transmission directions can utilize an ML-based encoder, we develop Active Feedback LightCode (ALC). [^4] We constrain the transmit power of the feedback encoder so that it matches the power constraint for the encoder transmitting the message, i.e., if $P=1$ and $T=9$, the RHS of (2) is $9$ for both the forward and feedback encoders. To implement ALC, we use the TWLC encoder, with User 2 always fed $\bm{b}_{2}=\bm{0}$ as a dummy initialization; the decoder is only fed $[\bm{y}_{2},{\color[rgb]{0,0,0}\bm{c}_{2}}]^{\intercal}$ since $\bm{b}_{2}$ provides no pertinent information. The overall loss is given by User 1’s CE loss. The layer dimensions of ALC in Tab. II are matched exactly to those in Figs. 12 & 13 in [12]; in Fig. 3(a), its dimensions are matched to TWLC’s. We note here that, under an appropriate training setup, it should in principle be possible to realize an implementation of ALC by training TWLC, disabling user $2$’s decoder, and sending only $\bm{0}$ from user $2$, thereby allowing the system to adaptively select its transmit mode based on the need of user $2$ to transmit bits; this, however, adds to the complexity and size of the deployed model.
TW GBAF-Based System (TWBAF): Beginning with the encoder/decoder modules presented in [9], we insert (with some abuse of notation) the function $\tanh{(\bm{X}-\bm{1}\mathbb{E}[\bm{X}]^{\intercal})}$ between the encoder’s output and the power reallocation mechanism, where $\bm{X}\in\mathbb{R}^{B\times\ell}$ is the raw encoder output, $B$ is the batch size, $\ell$ is the number of rows in the input matrix, and $\mathbb{E}[\bm{X}]\in\mathbb{R}^{\ell}$ is taken over the elements in the batch. This elementwise $\tanh$ function allows us to prevent divergent transmit symbols by mapping to the range $[-1,1]$ before power reweighting. Bits are not BPSK modulated at the model’s input. Without these updates, the TW model experienced training instability, failing to converge in all cases of $SNR_{1}=-1$ dB. Unlike [9], we only utilize time-indexed power reallocation via weights $\bm{w}_{i}=[w_{i}^{1},...,w_{i}^{T}]^{\intercal}$ rather than also reallocating along the block-dimension of each set of generated transmit symbols at time $t$; this permitted BLER improvement over reallocating along both dimensions. We use feature extractor C with ReLU activations, while matching the original specifications of the dimensions of the encoder/decoder attention modules from [9]. We forego the belief unit from [9] to keep the models comparably sized and reduce complexity. Despite having a similar FLOPS count to TWRNN (Sec. IV-A), TWBAF trained about 4×4$\times$4 × faster, which may be beneficial for on-device training.